Mr. Fragas business class has 96 students, classified by academic year and gender, as illustrated in the table. Mr. Fraga randomly chooses one 13 19 Step 2 of 2: What is the probability that he selects a junior, given that he chooses a male? Express your answer as a fraction or a decimal rounded to four decimal places How to enter your answer
Answer #1

From given table,

we get total number of males = 13 + 13 + 13 + 7 = 46

We want to find, the probability that the selects a junior, given that he chooses a male

P( junior | male ) = 13 / 46 = 0.2826

The probability is 0.2826